OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 1^2 + 4^2 + 7^2 + ... + (3n-2)^2 = n(6n^2-3n-1)/2 this is so hard, I need help :(

OpenStudy (anonymous):

@amistre64 do you know this by any chance??

OpenStudy (anonymous):

What part of the inductive process are you stuck at? Base Step? Inductive Step?

OpenStudy (anonymous):

ok I know that the first step is to have n=1 and when I did that it came out as true. But now I'm stuck on the S(k+1) step and on. It's just really confusing to me

OpenStudy (anonymous):

Alright, its good that you know the process. Your base step is done. Now you need to assume that:\[1^2+2^2+\cdots +k^2=\frac{1}{2}k(6k^2-3k-1)\]for some fixed natural number k. This is called your Inductive Hypothesis. What we need to prove (making sure to use the Inductive Hypothesis) is that:\[1^2+2^2+\cdots +k^2+(k+1)^2 =\frac{1}{2}(k+1)(6(k+1)^2-3(k+1)-1)\]

OpenStudy (anonymous):

Does what I'm doing make sense so far? (like how I got those equations, etc.)

OpenStudy (anonymous):

oh woah woah, my bad, I didn;t fully read the problem, let me correct some stuff.

OpenStudy (anonymous):

Inductive Hypothesis is:\[1^2+4^2+7^2+\cdots +(3k-2)^2=\frac{1}{2}k(6k^2-3k-1)\]

OpenStudy (anonymous):

What we want to prove is:\[1^2+4^2+7^2+\cdots +(3k-2)^2+(3(k+1)-2)^2\]\[=\frac{1}{2}(k+1)(6(k+1)^2-3(k+1)-1)\]

OpenStudy (anonymous):

Sry, I saw the squares and automatically assumed the problem was the sum of the first n squares.

OpenStudy (anonymous):

wait can u break this down for me I am still really confused! where did the k^2 and the 1/2k come from in the beginning?

OpenStudy (anonymous):

Ignore my earlier post, just look at the last three. Are there any questions where those equations came from?

OpenStudy (anonymous):

ok it's making more sense but just one question so far in 12+42+72+⋯+(3k−2)2+(3(k+1)−2)2 =12(k+1)(6(k+1)2−3(k+1)−1) where did u get the second +3(k+1)-2)^2

OpenStudy (anonymous):

That is the S(k+1) that you brought up earlier. The statement we are trying to prove is:\[1^2+4^2+7^2+\cdots +(3n-2)^2 =\frac{1}{2}n(6n^2-3n-1)\] This is S(n). S(k) is the inductive hypothesis, so I replace n with k. S(k+1) is what I want to prove, so I replace n with k+1.

OpenStudy (anonymous):

ahh okay I'm beginning to understand that part now, so what I'm wondering is since in the original we have (3n-2)^2 so if we were just replacing n with k wouldn't we just have (3k+1-2)^2? Because the way you did it you left it as but added (3k-2)^2+(3k+1-2)^2

OpenStudy (anonymous):

are you just suppose to keep it as (3k-2)^2 and then add it replacing it the second time?

OpenStudy (anonymous):

Ah, thats to help us actually solve the problem. To see why i left (3k-2)^2 in the S(k+1), lets look at some specific values of n, and see what the left side of the equation becomes. If n =1, then 3(1)-2=1, so the left side is:\[1^2\] If n=2, then 3(2)-2=4, so the left side is:\[1^2+4^2\] If n =3, then 3(3)-2=7, so the left side is:\[1^2+4^2+7^2\] Notice that each time you increase n by 1, all you're really doing is adding a new term to the end of the summation. That is why when I wrote down S(k+1), I left the (3k-2)^2 in the summation. Its still there, and (3(k+1)-2)^2 gets tacked on to the end.

OpenStudy (anonymous):

Its going to become super apparent why I left (3k-2)^2 in the summation when we start the inductive process.

OpenStudy (anonymous):

ahhh okay that makes way more sense now!! thank you for clearing it up :) what's the next step after setting that up? the inductive process?

OpenStudy (anonymous):

Right, now we actually have to prove that:\[1^2+4^2+7^2+\cdots +(3k-2)^2+(3(k+1)-2)^2\] \[=\frac{1}{2}(k+1)(6(k+1)^2-3(k+1)-1)\]

OpenStudy (anonymous):

But we have to make sure to use the Inductive Hyptothesis. That should be the first thing what we do. We are assuming that:\[1^2+4^2+7^2+\cdots +(3k-2)^2=\frac{1}{2}k(6k^2-3k-1)\]is true. Somehow we need to mathematically manipulate:\[1^2+4^2+7^2+\cdots +(3k-2)^2+(3(k+1)-2)^2\]and obtain:\[\frac{1}{2}(k+1)(6(k+1)^2-3(k+1)-1)\]

OpenStudy (anonymous):

This is the reason I left (3k-2)^2 in the summation. Note that:\[1^2+4^2+7^2+\cdots +(3k-2)^2+(3(k+1)-2)^2\]\[=\left(1^2+4^2+7^2+\cdots +(3k-2)^2\right)+(3(k+1)-2)^2\](all i did was group together all the terms but the last)

OpenStudy (anonymous):

By the inductive hypothesis, we know that:\[1^2+4^2+7^2+\cdots +(3k-2)^2=\frac{1}{2}k(6k^2-3k-1)\]So we can substitute this in for whats in the bigger parenthesis. Does this make sense?

OpenStudy (anonymous):

So performing this substitution gives us:\[\left(1^2+4^2+7^2+\cdots +(3k-2)^2\right)+(3(k+1)-2)^2=\]\[\left(\frac{1}{2}k(6k^2-3k-1)\right)+(3(k+1)-2)^2\]This step was really important, so if anything is fuzzy say so.

OpenStudy (anonymous):

wait okay. so I see that you substituted the 1/2k(6k^2−3k−1) in where the 1^2+4^2+7^2+⋯+(3k−2)^2 used to be, but why did you do that?? I'm sorry if that's a dumb question :( like I see what you're doing, I'm just a little bit unclear

OpenStudy (anonymous):

No no, questions are good, especially at this part. By substituting, I am using the inductive hypothesis. If you never use the inductive hypothesis, then you never did an inductive proof. Recall, we are trying to turn:\[1^2+4^2+7^2+\cdots +(3k-2)^2+(3(k+1)-2)^2\]into:\[\frac{1}{2}(k+1)(6(k+1)^2-3(k+1)-1)\]Now, without using the inductive hypothesis, that might be a pretty hard task (not impossible, but lengthy for sure). By using the inductive hypothesis, look at how much closer we are to our goal! We just need to formally manipulate:\[\frac{1}{2}k(6k^2-3k-1)+(3(k+1)-2)^2\]to reach our goal.

OpenStudy (anonymous):

Any more questions for anything up till now? The rest of the problem is algebra.

OpenStudy (anonymous):

Alright, to finish this problem, I'm going to fully expand both:\[\frac{1}{2}k(6k^2-3k-1)+(3(k+1)-2)^2\] and :\[\frac{1}{2}(k+1)(6(k+1)^2-3(k+1)-1)\]and show that they are equal. This will complete the proof by induction.

OpenStudy (anonymous):

Note that:\[\frac{1}{2}k(6k^2-3k-1)+(3(k+1)-2)^2\]\[=\frac{1}{2}(6k^3-3k^2-k)+(3k+1)^2\]\[=3k^3-\frac{3}{2}k^2-\frac{1}{2}k+9k^2+6k+1\]\[=3k^3+\frac{15}{2}k^2+\frac{11}{2}k+1\]All i did here was algebraic manipulation. Distributing, squaring, collecting like terms, etc.

OpenStudy (anonymous):

(If any of my steps are wrong let me know, i tend to make simple mistakes lol)

OpenStudy (anonymous):

Now Im going to fully expand:\[\frac{1}{2}(k+1)(6(k+1)^2-3(k+1)-1)\] and I hope to get the same thing as in the last post.

OpenStudy (anonymous):

Here we go (fingers crosses):\[\frac{1}{2}(k+1)(6(k+1)^2-3(k+1)-1)\]\[=\frac{1}{2}(6(k+1)^3-3(k+1)^2-(k+1))\]\[=\frac{1}{2}(6(k^3+3k^2+3k+1)-3(k^2+2k+1)-k-1)\]\[=\frac{1}{2}(6k^3+18k^2+18k+6-3k^2-6k-3-k-1)\]\[=\frac{1}{2}(6k^3+15k^2+11k+2)\]\[=3k^3+\frac{15}{2}k^2+\frac{11}{2}k+1\]

OpenStudy (anonymous):

So we did get the same thing, and we have (finally) proved that:\[1^2+4^2+7^2+\cdots +(3k-2)^2+(3(k+1)-2)^2\]is the same as:\[\frac{1}{2}(k+1)(6(k+1)^2-3(k+1)-1)\]This completes the inductive step of a proof by induction.

OpenStudy (anonymous):

Since we have completed both the base step and the inductive step, we have fully proven that:\[1^2+4^2+7^2+\cdots +(3n-2)^2=\frac{1}{2}n(6n^2-3n-1)\]for all natural numbers n.

OpenStudy (anonymous):

@joemath314159 Oh my gosh thank you so much!!! it really makes more sense now!